-3i-x+(2i)^4=3+2i/2-i
x=all complex numbers
Solve for x
Thank You in advance!!!

Answer by stanbon(75887)

-3i-x+(2i)^4=3+2i/2-i
---
-3i - x + 16 = 3 + i -i
-------
x = 13-3i
===================
Cheers,
Stan H.